# Improving Instruct Models for Free: A Study on Partial Adaptation

### v1

@@ Abstract
Instruct models, obtained from various instruction tuning or post-training steps, are commonly deemed superior and more usable than their base counterpart. While the model gains instruction following ability, instruction tuning may lead to forgetting the knowledge from pre-training or it may encourage the model being overly conversational or verbose. This, in turn, can lead to degradation of in-context few-shot learning performance. In this work, we study the performance trajectory between base and instruct models by scaling down the strength of instruction-tuning via the partial adaption method. We show that, across several model families and model sizes, reducing the strength of instruction-tuning results in material improvement on a few-shot in-context learning benchmark covering a variety of classic natural language tasks. This comes at the cost of losing some degree of instruction following ability as measured by AlpacaEval. Our study shines light on the potential trade-off between in-context learning and instruction following abilities that is worth considering in practice.
@@ 1 Introduction
However, both SFT and RLHF, to some degree, encourage the model to produce long and conversational responses. This may be an unwanted feature when testing on extractive and/or structured natural language processing (NLP) tasks such as classification, name entity recognition, or extractive question answering. In these cases, the responses need to be concise and exact, and any additional chattiness creates issues in parsing the responses. Before instruct models became available, this need was fulfilled decently by the emergent few-shot in-context learning (ICL) abilities of the base model Wei et al. (2022). Few previous studies touch on the pros and cons of base and instruct models. One example is Cuconasu et al. (2024) which shows how base models work better than instruct models on RAG-related tasks.
Our work aims to fill this gap and thoroughly explores the performance trajectory between base and instruct models. In order to study the learning dynamics between base and instruct models, we need access to the model checkpoints saved during instruct tuning, which are rarely available, especially for best performing open-weight models. Therefore as a surrogate of this Na et al. (2024), we resort to a simple training-free technique, partial adaptation (or PAd) Fleshman and Van Durme (2024), to scale the instruction-tuning strength in a post-hoc manner. Concretely, we create in-between models by partially adapting the base model (with weights $\mathbf{W_{B}}$) to instruct (with weights $\mathbf{W}_{I}$): $M_{\lambda}$ with weights $\mathbf{W_{B}}+\lambda\mathbf{A}\;(\lambda\in[0,1])$ where $\mathbf{A}\equiv\mathbf{W}_{I}-\mathbf{W}_{B}$. Hence, $M_{0}$ is the base model and $M_{1}$ is the instruct model (see Section 2 for more details).
@@
In summary, through this comprehensive analysis, we demonstrate that the best ICL model is not necessarily the instruct model. We believe partial adaptation represents a training-free yet effective option worth exploring when dealing with ICL tasks that are structured, more extractive in nature, or requiring shorter answers. We hope our study highlights the opportunities and can inspire future work in better understanding the learning dynamics in LLM post training.
@@ 2 Preliminary: Partial Adaptation
The partially adapted model can also be viewed as the weighted average between base and instruct models. Hence, we consider a new model $M_{\lambda}$ with weights $(1-\lambda)\mathbf{W}_{B}+\lambda\mathbf{W}_{I}$, so that $M_{0}$ and $M_{1}$ correspond to the base and instruct models respectively. Open-weight models that we consider are listed in Table  1.111For all of the models, except Mixtral 8$\times$22B, the embedding lookup tables of the base and instruct versions are aligned, so merging is straightforward. For Mixtral 8$\times$22B, there are additional special tokens in the vocabulary of the instruct model. We take care of this by applying $\lambda=1$ for those weights that are only present in the instruct model. In practice, we enumerate $\lambda$ from $\{0,\frac{1}{8},\frac{2}{8},\frac{3}{8},\frac{4}{8},\frac{5}{8},\frac{6}{8},\frac{7}{8},1\}$.
@@ 3.1 In-Context Learning Benchmark
Our primary goal is to measure performance on few-shot in-context learning. We assemble a benchmark of various classic NLP tasks to test a variety of natural language abilities. The composition of the benchmark is shown in Table 2 and described in details in Appendix A.1. We particularly include tasks from the financial domain because classic structured NLP tasks (classification, name entity recognition, extractive QA) widely appear in financial data analysis. Each dataset is tested in a few-shot manner, where the number of shots is displayed in Table 2. Shot selection is random and done independently for each example.
@@
When a single dataset is evaluated in multiple ways (different prompts or different evaluation styles: MC vs. FMC vs. G), we aggregate these individual scores by taking their maximum. The final score that we use to rank the various models is the average of these aggregated dataset-level scores. More details about the templates and metrics that we use in our evaluation protocol are presented in Appendix A.3 and A.4.
@@ 3.2 AlpacaEval
Instruction following is a broad concept. In this work, we refer to it as the model’s ability to answer open-ended questions from users, as exemplified by Chatbot Arena Chiang et al. (2024). [^3] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard Here, we test on AlpacaEval 2.0 Dubois et al. (2024), which has a Spearman correlation of 0.98 with Chatbot Arena while being cost-efficient. For each value of $\lambda$, we obtain the length-controlled win-rate of $M_{\lambda}$ against GPT-4 Preview (11/06) Li et al. (2023) judged by GPT-4o. [^4] The GPT-4o version that we use is the May 2024 one.
@@
Llama-2 7B 51.9/50.5 52.85/8 --4.35
Llama-2 70B 64.8/60.9 65.92/8 --16.64
Llama-3 8B 59.4/58.3 61.94/8 --15.81
Llama-3 70B 68.5/66.6 70.43/8 --6.02
Llama-3.1 8B 59.3/61.2 62.45/8 --5.58
Llama-3.1 70B 69.0/69.8 71.34/8 --5.30
Llama-3.2 1B 43.2/45.4 45.95/8 --8.93
Llama-3.2 3B 53.6/55.6 57.25/8 --8.89
Llama-3.3 70B 69.0/70.0 71.45/8 --0.93
Mistral 7B v0.1 56.6/53.3 58.62/8 --6.73
Mistral 7B v0.3 57.1/58.9 59.56/8 --1.57
Mistral Nemo 12B 62.5/63.1 64.15/8 --5.70
Mixtral 8x7B v0.1 62.2/61.4 63.23/8 --14.48
@@
Gemma-2 9B 57.6/58.2 59.64/8 --6.52
OLMo 7B 0724 51.1/49.1 52.75/8 --6.79
OLMo 2 7B 1124 55.7/55.4 57.94/8 --9.95
OLMo 2 13B 1124 60.2/61.1 61.56/8 --4.43
Table 1: For each of the models (LLama-2 Touvron et al. (2023), Llama-3 Dubey et al. (2024); Meta (July 2024, September 2024, December 2024), Mistral Jiang et al. (2023), Mistral-NeMo mistral.ai (July 2024), Mixtral Jiang et al. (2024); mistral.ai (April 2024), Gemma-2 Riviere et al. (2024), OLMo Groeneveld et al. (2024), and OLMo-2 OLMo et al. (2024)) in the first column we report the base and instruct baseline performance on the benchmark, together with the best performance obtained by varying $\lambda$ and the best value $\lambda^{*}$ at which peak performance is achieved. The last columns reports the absolute change in win rate for the best PAd model with respect to the instruct version as determined by AlpacaEval 2.0. NA is because $\lambda^{*}=0$ and we don’t evaluate AlpacaEval on the base model when chat template does not exist.
Figure 2: Performance on AlpacaEval 2.0: fractional difference (percent value) between the length controlled win rate of each partially adapted model $M_{\lambda}$ and the instruct baseline $M_{1}$ against GPT-4 Preview (11/06).
@@ 4 Results
Figure 1 and Figure 2 illustrate the relative performance change of each partially adapted model $M_{\lambda}$ against the instruct model $M_{1}$ on ICL and and AlpacaEval benchmark, respectively. And Figure 4 and Figure 3 in Appendix B shows the corresponding absolute values. We summarize the absolute performance of base/instruct models and the best partially adapted models as well as the best $\lambda^{*}$ in Table 1.
@@
The improvement on ICL is at the cost of losing some instruction following abilities as measured by the AlpacaEval 2.0 win rate shown in Figure 2 and the last column of Table 1. In Table 1, $\delta_{\textrm{wr}}\equiv{\textrm{wr}}_{M_{\lambda^{*}}}-{\textrm{wr}}_{M_{1}}$ represents the absolute difference in win rate between the best PAd model for ICL ($M_{\lambda^{*}}$) and the instruct version ($M_{1}$). As shown in Figure 2, the best win rate is mostly achieved by the instruct model, except for a few cases where a marginally higher win rate is achieved when $10.6<\lambda<$ 1.
ICL can be improved with a small drop of instruction following abilities. We notice that for many models, especially the larger ones, the win-rate curve saturates to the instruct value for $\lambda$ values well below 1. This implies that there are values of $\lambda$, in the range $\lambda^{*}\leq\lambda<1$, where the AlpacaEval 2.0 performance does not drop significantly, yet there is still a gain on the ICL benchmark due to PAd. For instance, by allowing at most a $1\%$ relative win rate decrease from the instruct model on AlpacaEval 2.0, we can get a $+5.9\%$ relative improvement on the ICL benchmark performance for Llama-2 70B ($\lambda=0.625$), $+4.9\%$ for Llama-3 70B ($\lambda=0.625$), $+1.7\%$ for Llama-3.3 70B ($\lambda=0.75$).
@@ 5 Conclusion and Future Work
Nonetheless, the observation that instruction following performance for larger models is not very sensitive to $\lambda$ when $\lambda\lesssim 1$ suggests that scaling down instruction tuning strength to a small degree would consistently be beneficial. Hence, it would make sense to apply PAd at the end of post-training (e.g., replacing $M_{1}$ with $M_{\lambda^{*}}$) to further boost model performance. This might have already happened as Llama 3.3 Meta (December 2024) used an annealing technique to average model checkpoints, and we also observed that PAd boosts Llama 2 ICL performance much more than Llama 3.3.
Future work can focus on better understanding why PAd improves ICL performance by studying its impact on each stage of supervised fine-tuning or RL. Another avenue of investigation is a thorough comparison of the training dynamics during instruction tuning with the model trajectory defined by varying $\lambda$ in PAd. It has been suggested that the latter may indeed recapitulate the full training dynamics Na et al. (2024).
@@
Our method is evaluated on a collection of 21 common datasets used for in-context learning spanning 6 broad types of tasks. The collection may however not be fully representative of the model performance or its performance on other specific tasks. Further, we limit our study to models primarily trained on English data and tasks in English, hence we did not test the generalizability to other languages and multi-lingual models, and leave this to future work.
@@ Acknowledgements
We thank Steven Lu and Shijie Wu for their involvement in the development of the in-context learning benchmark.
@@ A.1 Datasets
  • Language processing and understanding: we include five classic language processing or understanding tasks. BBH (NLP) are NLP tasks from Big Bench Hard Suzgun et al. (2023), e.g., movie recommendation. FiQA (SA) and FFB (SA) are two sentiment analysis tasks, Headline is a headline classification tasl, Flue (NER) is a name entity recognition task, and all these four datasets are from FLUE (Financial Language Understanding Evaluation) benchmark Shah et al. (2022).
@@ A.2 Evaluation Tasks
    $N(c)$ is a possibly choice dependent normalization that we use to ameliorate possible biases of the model likelihood Zheng et al. (2023). We consider 3 possibilities for $N$
@@ A.3 Templates
1{set ENUM = ABCDEFGHIJKLM %}The following are multiple choice questions (with answers) about {{subject}}.
2{for example in examples %}
@@
4{for choice in example.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
5{endfor %}Answer: ({{ ENUM[example.gold] }}) {{ example.choices[example.gold] }}
6{endfor %}
@@
8{for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{endfor %}Answer:
@@
1{for example in examples %}{{ example.question }}
@@
4{for choice in ex.choices %}- {{ choice }}
5{endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
10{for choice in choices %}- {{ choice }}
11{endfor %}|||Answer:
@@
1{set ENUM = ABCDEFGHIJKLM %}{for ex in examples %}{{ ex.sentence }}
@@
3{for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
4{endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
8{for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{endfor %}Answer:
@@
1{for example in examples %}{{ example.question }}
@@
1{for example in examples %}{{ example.question }}
@@
3{for choice in ex.choices %}- {{ choice }}
4{endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
8{for choice in choices %}- {{ choice }}
9{endfor %}|||Answer:
@@
1{set ENUM = ABCDEFGHIJKLM %}{for ex in examples %}{{ ex.question }}
2{for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
3{endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
5{endfor %}{{ question }}
6{for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
7{endfor %}Answer:
@@
4{for choice in ex.choices %}- {{ choice }}
5{endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
10{for choice in choices %}- {{ choice }}
11{endfor %}|||Answer:
@@
1{set ENUM = ABCDEFGHIJKLM %}{for ex in examples %}{{ ex.sentence }}
@@
3{for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
4{endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
8{for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{endfor %}Answer:
@@
1{for example in examples %}{{ example.question }}
@@
1{for example in examples %}{{ example.question }}{{ example.choices[example.gold] }}

### v2

@@ Abstract
Instruct models, obtained from various instruction tuning or post-training steps, are commonly deemed superior and more usable than their base counterpart. While the model gains instruction following ability, instruction tuning may lead to forgetting the knowledge from pre-training or it may encourage the model to become overly conversational or verbose. This, in turn, can lead to degradation of in-context few-shot learning performance. In this work, we study the performance trajectory between base and instruct models by scaling down the strength of instruction-tuning via the partial adaption method. We show that, across several model families and model sizes, reducing the strength of instruction-tuning results in material improvement on a few-shot in-context learning benchmark covering a variety of classic natural language tasks. This comes at the cost of losing some degree of instruction following ability as measured by AlpacaEval. Our study shines light on the potential trade-off between in-context learning and instruction following abilities that is worth considering in practice.
@@ 1 Introduction
However, both SFT and RLHF, to some degree, encourage the model to produce long and conversational responses. This may be an unwanted feature when testing on extractive and/or structured natural language processing (NLP) tasks such as classification, named entity recognition, or extractive question answering. In these cases, the responses need to be concise and exact, and any additional chattiness creates issues in parsing the responses. Before instruct models became available, this need was fulfilled reasonably well by the emergent few-shot in-context learning (ICL) abilities of the base model Wei et al. (2022). Few previous studies touch on the pros and cons of base and instruct models. One example is Cuconasu et al. (2024) which shows how base models work better than instruct models on RAG-related tasks.
Our work aims to fill this gap and thoroughly explores the performance trajectory between base and instruct models. In order to study the learning dynamics between base and instruct models, we need access to the model checkpoints saved during instruct tuning, which are rarely available, especially for best performing open-weight models. Therefore as a surrogate of this Na et al. (2024), we resort to a simple training-free technique, partial adaptation (or PAd) Fleshman and Van Durme (2024), to scale the instruction-tuning strength in a post-hoc manner. Specifically, we create in-between models by partially adapting the base model (with weights $\mathbf{W_{B}}$) to instruct (with weights $\mathbf{W}_{I}$): $M_{\lambda}$ with weights $\mathbf{W_{B}}+\lambda\mathbf{A}\;(\lambda\in[0,1])$ where $\mathbf{A}\equiv\mathbf{W}_{I}-\mathbf{W}_{B}$. Hence, $M_{0}$ is the base model and $M_{1}$ is the instruct model (see Section 2 for more details).
@@
In summary, through this comprehensive analysis, we demonstrate that the best ICL model is not necessarily the instruct model. We believe partial adaptation represents a training-free yet effective option worth exploring when dealing with ICL tasks that are structured, more extractive in nature, or requiring shorter answers. We hope our study highlights the opportunities and can inspire future work in better understanding the learning dynamics in LLM post-training.
@@ 2 Preliminary: Partial Adaptation
[…]
@@ 3.1 In-Context Learning Benchmark
Our primary goal is to measure performance on few-shot in-context learning. We assemble a benchmark of various classic NLP tasks to test a variety of natural language abilities. The composition of the benchmark is shown in Table 2 and described in detail in Appendix A.1. We particularly include tasks from the financial domain because classic structured NLP tasks (classification, named entity recognition, extractive QA) are common in financial data analysis. Each dataset is tested in a few-shot manner, where the number of shots is displayed in Table 2. Shot selection is random and done independently for each example.
@@
When a single dataset is evaluated in multiple ways (different prompts or different evaluation styles: MC vs. FMC vs. G), we aggregate these individual scores by taking their maximum. All metric sores are in a scale of 0 to 100. Therefore, we are able to average dataset-level scores into one single model-level score. More details about the templates and metrics that we use in our evaluation protocol are presented in Appendix A.3 and A.4.
@@ 3.2 AlpacaEval
Instruction following is a broad concept. In this work, we refer to it as the model’s ability to answer open-ended questions from users, as exemplified by Chatbot Arena Chiang et al. (2024). Here, we test on AlpacaEval 2.0 Dubois et al. (2024), which has a Spearman correlation of 0.98 with Chatbot Arena while being cost-efficient. For each value of $\lambda$, we obtain the length-controlled win-rate of $M_{\lambda}$ against GPT-4 Preview (11/06) Li et al. (2023) judged by GPT-4o. [^3] The GPT-4o version that we use is the May 2024 one.
@@
Llama-2 7B 51.9/50.5 52.85/8 -4.35
Llama-2 70B 64.8/60.9 65.92/8 -16.64
Llama-3 8B 59.4/58.3 61.94/8 -15.81
Llama-3 70B 68.5/66.6 70.43/8 -6.02
Llama-3.1 8B 59.3/61.2 62.45/8 -5.58
Llama-3.1 70B 69.0/69.8 71.34/8 -5.30
Llama-3.2 1B 43.2/45.4 45.95/8 -8.93
Llama-3.2 3B 53.6/55.6 57.25/8 -8.89
Llama-3.3 70B 69.0/70.0 71.45/8 -0.93
Mistral 7B v0.1 56.6/53.3 58.62/8 -6.73
Mistral 7B v0.3 57.1/58.9 59.56/8 -1.57
Mistral Nemo 12B 62.5/63.1 64.15/8 -5.70
Mixtral 8x7B v0.1 62.2/61.4 63.23/8 -14.48
@@
Gemma-2 9B 57.6/58.2 59.64/8 -6.52
OLMo 7B 0724 51.1/49.1 52.75/8 -6.79
OLMo 2 7B 1124 55.7/55.4 57.94/8 -7.41
OLMo 2 13B 1124 60.2/61.1 61.56/8 -3.98
Table 1: For each of the models (LLama-2 Touvron et al. (2023), Llama-3 Dubey et al. (2024); Meta (2024, September 2024, December 2024), Mistral Jiang et al. (2023), Mistral-NeMo mistral.ai (July 2024), Mixtral Jiang et al. (2024); mistral.ai (April 2024), Gemma-2 Riviere et al. (2024), OLMo Groeneveld et al. (2024), and OLMo-2 OLMo et al. (2024)) in the first column we report the base and instruct baseline performance on the benchmark, together with the best performance obtained by varying $\lambda$ and the best value $\lambda^{*}$ at which peak performance is achieved. The last column reports the absolute change in win rate for the best PAd model with respect to the instruct version as determined by AlpacaEval 2.0. NA is because $\lambda^{*}=0$ and we don’t evaluate AlpacaEval on the base model when chat template does not exist.
Figure 2: Performance on AlpacaEval 2.0: fractional difference (percent value) between the length-controlled win rate of each partially adapted model $M_{\lambda}$ and the instruct baseline $M_{1}$ against GPT-4 Preview (11/06).
@@ 4 Results
Figure 1 and Figure 2 illustrate the relative performance change of each partially adapted model $M_{\lambda}$ against the instruct model $M_{1}$ on ICL and AlpacaEval benchmark, respectively. And Figure 4 and Figure 3 in Appendix B shows the corresponding absolute values. We summarize the absolute performance of base/instruct models and the best partially adapted models as well as the best $\lambda^{*}$ in Table 1.
@@
The improvement on ICL is at the cost of losing some instruction following abilities as measured by the AlpacaEval 2.0 win rate shown in Figure 2 and the last column of Table 1. In Table 1, $\delta_{\textrm{wr}}\equiv{\textrm{wr}}_{M_{\lambda^{*}}}-{\textrm{wr}}_{M_{1}}$ represents the absolute difference in win rate between the best PAd model for ICL ($M_{\lambda^{*}}$) and the instruct version ($M_{1}$). As shown in Figure 2, the best win rate is mostly achieved by the instruct model, except for a few cases where a marginally higher win rate is achieved when $0.6<\lambda<1$.
ICL can be improved with a small drop of instruction following abilities. We notice that for many models, especially the larger ones, the win-rate curve saturates to the instruct value for $\lambda$ values well below 1. This implies that there are values of $\lambda$, in the range $\lambda^{*}\leq\lambda<1$, where the AlpacaEval 2.0 performance does not drop significantly, yet there is still a gain on the ICL benchmark due to PAd. For instance, by allowing a relative win-rate decrease of at most 1% from the instruct model on AlpacaEval 2.0, we can get a $5.9\%$ relative improvement on the ICL benchmark performance for Llama-2 70B ($\lambda=0.625$), $4.9\%$ for Llama-3 70B ($\lambda=0.625$), $1.7\%$ for Llama-3.3 70B ($\lambda=0.75$).
@@ 5 Conclusion and Future Work
Nonetheless, the observation that instruction following performance for larger models is not very sensitive to $\lambda$ when $\lambda\lesssim 1$ suggests that slightly scaling down instruction tuning strength would consistently be beneficial. Hence, it would make sense to apply PAd at the end of post-training (e.g., replacing $M_{1}$ with $M_{\lambda^{*}}$) to further boost model performance. This might have already happened as Llama 3.3 Meta (December 2024) used an annealing technique to average model checkpoints, and we also observed that PAd boosts Llama 2 ICL performance much more than Llama 3.3.
Future work can focus on better understanding why PAd improves ICL performance by studying its impact on each stage of supervised fine-tuning or RL. Another avenue of investigation is a thorough comparison of the training dynamics during instruction tuning with the model trajectory defined by varying $\lambda$ in PAd. It has been suggested that the latter may indeed replicate the full training dynamics Na et al. (2024).
@@
Our in-context learning benchmark is a collection of 21 common datasets spanning 6 broad types of tasks. The collection may however not be fully representative of the model ICL performance or its performance on other specific tasks. Similarly, we benchmark instruction following ability on AlpacaEval 2.0 Dubois et al. (2024), which has a Spearman correlation of 0.98 with Chatbot Arena. However, it may not be fully representative of the model true instruction following performance. Further, we limit our study to models primarily trained on English data and tasks in English, hence we leave testing the generalizability to other languages and multilingual models to future work.
@@ Acknowledgements
We thank Steven Lu and Shijie Wu for their contributions to developing the in-context learning benchmark.
@@ A.1 Datasets
  • Language processing and understanding: we include five classic language processing or understanding tasks. BBH (NLP) are NLP tasks from Big Bench Hard Suzgun et al. (2023), e.g., movie recommendation. FiQA (SA) and FFB (SA) are two sentiment analysis tasks, Headline is a headline classification task, Flue (NER) is a named entity recognition task, and all these four datasets are from FLUE (Financial Language Understanding Evaluation) benchmark Shah et al. (2022).
@@ A.2 Evaluation Tasks
    $N(c)$ is a possibly choice dependent normalization that we use to ameliorate possible biases of the model likelihood Zheng et al. (2023). We consider 3 possibilities for NN
@@ A.3 Templates
1{% set ENUM = ABCDEFGHIJKLM %}The following are multiple choice questions (with answers) about {{subject}}.
2{% for example in examples %}
@@
4{% for choice in example.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
5{% endfor %}Answer: ({{ ENUM[example.gold] }}) {{ example.choices[example.gold] }}
6{% endfor %}
@@
8{% for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{% endfor %}Answer:
@@
1{% for example in examples %}{{ example.question }}
@@
4{% for choice in ex.choices %}- {{ choice }}
5{% endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
10{% for choice in choices %}- {{ choice }}
11{% endfor %}|||Answer:
@@
1{% set ENUM = ABCDEFGHIJKLM %}{% for ex in examples %}{{ ex.sentence }}
@@
3{% for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
4{% endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
8{% for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{% endfor %}Answer:
@@
1{% for example in examples %}{{ example.question }}
@@
1{% for example in examples %}{{ example.question }}
@@
3{% for choice in ex.choices %}- {{ choice }}
4{% endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
8{% for choice in choices %}- {{ choice }}
9{% endfor %}|||Answer:
@@
1{% set ENUM = ABCDEFGHIJKLM %}{% for ex in examples %}{{ ex.question }}
2{% for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
3{% endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
5{% endfor %}{{ question }}
6{% for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
7{% endfor %}Answer:
@@
4{% for choice in ex.choices %}- {{ choice }}
5{% endfor %}Answer: {{ ex.choices[ex.gold] }}
@@
10{% for choice in choices %}- {{ choice }}
11{% endfor %}|||Answer:
@@
1{% set ENUM = ABCDEFGHIJKLM %}{% for ex in examples %}{{ ex.sentence }}
@@
3{% for choice in ex.choices %}({{ ENUM[loop.index0] }}) {{ choice }}
4{% endfor %}Answer: ({{ ENUM[ex.gold] }}) {{ ex.choices[ex.gold] }}
@@
8{% for choice in choices %}({{ ENUM[loop.index0] }}) {{ choice }}
9{% endfor %}Answer:
@@
1{% for example in examples %}{{ example.question }}
@@
1{% for example in examples %}{{ example.question }}{{ example.choices[example.gold] }}